OpenStudy (anonymous):

i'm in algebra 1 barely. & I need help with this problem. given that, 2(x-1)=3(5+x)-12, what is the value of 3x to the second power+2?

OpenStudy (anonymous):

First you'll have to look for the value of x So continuing from your equation, \[2x-2=15+3x-12\]\[x=-5\] Then substitute your value of x into your given question which is \[3x^{2}+2\]\[=3(-5)^{2}+2\]\[=77\]

OpenStudy (anonymous):

that's not what I got. I got x=5.

OpenStudy (anonymous):

hmmm i too made a mistake it is 3x^2+2 i took (3x)^(+2)

OpenStudy (anonymous):

next time to help make sense of your problem use parenthesis. I was under the impression it was x^2 +another 2 making it 3x^4. But first answer is right

OpenStudy (anonymous):

@alyssa: You'll get the same result which is 77 cause square of 5 and square of -5 are both 25. @sheg: Really? Cause that wasn't how we usually pronounce it so perhaps I'm wrong. :P

OpenStudy (anonymous):

multiply out using the distributive law \[2x-2=15+3x-12\] combine like terms \[2x-2=3+3x\] subtract 2x from both sides \[-2=3+x\] subtract 3 from both sides \[-5=x\]

OpenStudy (anonymous):

now if it is \[3x^2\] you get \[3\times (-5)^2=3\times 25=75\]

OpenStudy (anonymous):

and if it is \[3x^2+2\] you get \[75+2=77\] but if it is \[(3x)^2+2\] then you get something else

OpenStudy (anonymous):

wait, what does " ^ " stand for? o:

OpenStudy (anonymous):

is the problem \[3x^2\]?

OpenStudy (anonymous):

3x^2 means \[3x^2\]

OpenStudy (anonymous):

help me earn medals? :D and thanks for your help. but I still want to know what "^" stands for.

OpenStudy (anonymous):

We write that as it's how we key in the power of a number into the calculator